# Mexa: Multilingual Evaluation of English-Centric LLMs via Cross-Lingual Alignment

### v1

@@ 4.1 Parallel Data
We calculate the Mexa score using the parallel datasets of FLORES-200 (NLLB Team et al., 2022) and the Bible (Mayer & Cysouw, 2014). While there are other high-quality parallel datasets, such as NTREX-128 (Federmann et al., 2022), IN22 (Gala et al., 2023), OPUS-100 (Zhang et al., 2020), Europarl (Koehn, 2005), OpenSubtitles (Lison & Tiedemann, 2016), TED2020 (Reimers & Gurevych, 2020), and Tatoeba (Tatoeba Community, 2006), we specifically chose FLORES-200 due to its high quality and support for a wide range of languages, while the Bible dataset was selected for its extensive language coverage.
FLORES-200 is a parallel corpus, where the English subset is sourced from Wikimedia projects. Each English sentence has been translated into 204 distinct language-script combinations, these translations have been verified by humans. The dataset contains 997 sentences for development, 1012 sentences for dev-test, and 992 sentences for testing. As the FLORES-200 test set is not publicly accessible, we use the dev-test set as our FLORES parallel test corpus, in line with previous studies. For faster computation, we only consider the first 100 sentences from each language. As shown in Appendix A.4, the odds of the Mexa score randomly achieving a high value with 100 sentences are very slim.
The Parallel Bible (Mayer & Cysouw, 2014) covers a very large number of languages. From this resource, we managed to create a subcorpus, a super parallel dataset of the Bible, with 1,401 language-script labels, each containing 103 sentences (i.e., Bible verses).111hf.co/datasets/cis-lmu/sPBC This corpus includes many low-resource languages, many of which are not covered by existing language technologies (Joshi et al., 2020), and Mexa can be adopted since only parallel data is needed. We use all the 103 sentences from each language.
@@ 4.2 Models
For our experiments, we select models with around 7B parameters, which are considered a base size in the LLM community. The state-of-the-art open-weight models in this range include Llama 1, 2, 3, and 3.1 (Touvron et al., 2023a; b; Meta, 2024; Dubey et al., 2024), Gemma 1 and 2 (Gemma Team et al., 2024a; b), Mistral 0.3 (Jiang et al., 2023a), and the open-source model OLMo 1.7 (Groeneveld et al., 2024). We also select a larger model, Llama 3.1 70B, to show that our findings hold even when scaling further. To apply Mexa, we need to access model weights to compute input sentence embeddings for each layer. We use three popular open-weight model families: Llama, Gemma, and Mistral. As a less multilingual version of state-of-the-art LLMs, we include OLMo, which is trained on a more English-centric corpus of Dolma (Soldaini et al., 2024). Although there are multilingual models such as PolyLM (with support of 18 languages), XGLM (Lin et al., 2022) (with support of 30 languages) and BLOOM (BigScience Workshop et al., 2023) (with support of 46 languages) our focus here is on LLMs which are state-of-the-art in English based tasks such as MMLU (Stanford CRFM, 2024).
4.3 Benchmarks
Among the existing evaluation benchmarks in Table 4 from Appendix A.2, we chose the Belebele benchmark (Bandarkar et al., 2024), m-ARC (Lai et al., 2023b), and m-MMLU (Lai et al., 2023b), which support the highest number of high-, medium-, and low-resource languages and are directly related to natural understanding tasks, which is the primary focus of this paper.
Belebele is a multiple-choice reading comprehension task designed to assess language models across a range of high-, medium-, and low-resource languages. Each question in the dataset is accompanied by four possible answers and is linked to a brief passage from the FLORES-200 dataset (NLLB Team et al., 2022). The human annotation process was carefully curated to generate questions that effectively differentiate between various levels of language comprehension, supported by rigorous quality checks. Belebele supports 122 distinct labels (language-script combinations) corresponding to 115 distinct languages. However, FLORES-200 does not support 5 of these labels, corresponding to Romanized versions of 5 Indic languages. Therefore, we conducted our analysis between the FLORES-200 and Belebele benchmarks on 117 common labels. Additionally, there are 102 common labels between the Bible parallel data and Belebele benchmark.
Both ARC Challenge (Clark et al., 2018) and MMLU (Hendrycks et al., 2021) are also set up as multiple-choice question-answering tasks, but they focus on different types of knowledge and reasoning skills. ARC Challenge is classified as a common-sense reasoning task, consisting of grade-school level science questions, while MMLU consists of questions across a wide range of subjects, including humanities, social sciences, and more. Lai et al. (2023b) used GPT-3.5-turbo (OpenAI, 2022) and a translation prompt to translate examples from both datasets and create m-ARC and m-MMLU in 31 languages (excluding English). Later, m-MMLU was expanded to also include Icelandic (isl_Latn) and Norwegian (nob_Latn). The Icelandic portion was translated using the Mideind.is, while Norwegian was generated with DeepL.com.222hf.co/datasets/alexandrainst/m_mmlu m-MMLU consists of 277 questions in its training set, 13,258 in the test set and 1,433 in the validation set. m-ARC consists of 1,116 questions in the training set, 1,169 in the test set, and 298 in the validation set. We use the entire test set for each of these benchmarks to evaluate the models, except in one case. For Llama 3.1 70B, we use the first 500 questions of m-MMLU instead of the whole set due to resource constraints. Since the selected LLMs used in our experiment are not instruction-tuned, we use 5-shot in-context learning with the lm-evaluation-harness framework, employing log-likelihood-based multiple-choice scoring. Other settings, such as prompt templates, are configured according to the framework’s default (Gao et al., 2023; Biderman et al., 2024).
4.4 Evaluation Measures
We use Pearson correlation to assess the strength of the correlation between Mexa and downstream performance on our evaluation benchmarks. Pearson correlation is a statistical measure that calculates the strength and direction of the linear relationship between two variables. A high Pearson correlation would indicate that Mexa provides a reliable assessment of multilingual capabilities in English-centric LLMs.
5 Results
@@
| Belebele (avg., L = 116) | 0.7093 | 0.5633 | 0.7684 | 0.5705 | 0.5533 | 0.3028 | 0.2755 | 0.4457 | 0.3627 | 0.5057 | |
| m-MMLU (avg., L = 33) | 0.5582 | 0.4734 | 0.6384 | 0.4720 | 0.4664 | 0.3260 | 0.2807 | 0.4207 | 0.3390 | 0.4416 | |
| m-ARC (avg., L = 31) | 0.4779 | 0.4220 | 0.5054 | 0.3941 | 0.3892 | 0.3174 | 0.2970 | 0.3662 | 0.2731 | 0.3825 | |
| FLORES | (avg., L = 116) | 0.5088 | 0.3815 | 0.4110 | 0.3963 | 0.3939 | 0.0866 | 0.1946 | 0.2642 | 0.0413 | 0.2976 |
| (avg., L = 116) | 0.7194 | 0.5872 | 0.7725 | 0.6538 | 0.6520 | 0.2464 | 0.3579 | 0.4716 | 0.1965 | 0.5175 | |
| Bible | (avg., L = 101) | 0.3568 | 0.2152 | 0.3169 | 0.2103 | 0.2026 | 0.1246 | 0.0908 | 0.1198 | 0.0121 | 0.1832 |
| (avg., L = 101) | 0.6076 | 0.4021 | 0.6599 | 0.4212 | 0.4190 | 0.2724 | 0.2357 | 0.2606 | 0.0319 | 0.3678 |
Table 1 presents the downstream performance of the selected models across three benchmarks, along with Mexa scores from two parallel datasets. Notably, among models with parameter sizes ranging from 7 to 9 billion, both Gemma 2 and Llama 3.1 outperform the others in terms of non-English downstream performance and Mexa scores. The Llama 3.1 and Llama 3 models exhibit similar alignment and downstream task performance; yet, both represent substantial advancements compared to Llama 2. Moreover, results for the Llama 3.1-70B model indicate that scaling can significantly enhance alignment when compared to its smaller version. Interestingly, while Mistral achieves comparable results to Gemma 1 on English benchmarks, it demonstrates inferior alignment, which likely accounts for its reduced performance on non-English tasks. Furthermore, the Llama 2 model achieves higher Mexa scores than OLMo, indicating better alignment. However, due to Llama 2’s weaker performance on English tasks, it fails to transfer this alignment effectively, leading to comparable non-English performance between Llama 2 and OLMo. This observation is further explored in Section 5.2, where we normalize the expected performance based on the pivot language, namely English.

### v2

@@ 4.1 Parallel Data
We calculate the MEXA score using the parallel datasets of FLORES-200 (NLLB Team et al., 2022) and the Bible (Mayer and Cysouw, 2014). While there are other high-quality parallel datasets, such as NTREX-128 (Federmann et al., 2022), IN22 (Gala et al., 2023), OPUS-100 (Zhang et al., 2020), Europarl (Koehn, 2005), OpenSubtitles (Lison and Tiedemann, 2016), TED2020 (Reimers and Gurevych, 2020), and Tatoeba (Tatoeba Community, 2006), we chose FLORES-200 due to its high quality and support for a wide range of languages, and the Bible dataset was selected for its extensive language coverage.
FLORES-200 is a parallel corpus with English sentences from Wikimedia translated into 204 language-script pairs, verified by humans. It includes 997 dev, 1012 dev-test, and 992 test sentences. As the test set isn’t public, we use the dev-test set as our test corpus, following prior work. For faster computation, we consider only the first 100 sentences from each language. As shown in Section 3.2, this is sufficient to ensure MEXA’s robustness, as the odds of the MEXA score randomly achieving a high value with 100 sentences are very slim. This choice also enables scaling to more languages, many of which lack enough parallel samples.
The Parallel Bible (Mayer and Cysouw, 2014) covers a very large number of languages. From this resource, we managed to create a subcorpus, a super parallel dataset of the Bible, with 1,401 language-script labels, each containing 103 sentences (i.e., Bible verses).111hf.co/datasets/cis-lmu/sPBC This corpus includes many low-resource languages, many of which are not covered by existing language technologies (Joshi et al., 2020), and MEXA can be adopted since only parallel data is needed. We use all 103 sentences from each language.
@@ 4.2 Models
For our experiments, we select models with around 7B parameters, which are considered a base size in the LLM community. The state-of-the-art open-weight models in this range, as measured by performance on English-based tasks such as MMLU (Stanford CRFM, 2024), include Llama 1, 2, 3, and 3.1 (Touvron et al., 2023a, b; Dubey et al., 2024), Gemma 1 and 2 (Gemma Team et al., 2024a, b), Mistral 0.3 (Jiang et al., 2023a), and the open-source model OLMo 1.7 (Groeneveld et al., 2024). We also select a larger model, Llama 3.1 70B, to show that our findings hold even when scaling up further. To apply MEXA, we need to access model weights to compute input sentence embeddings for each layer. We use three popular open-weight model families: Llama, Gemma, and Mistral. As a less multilingual version of state-of-the-art LLMs, we include OLMo, which is trained on a more English-centric corpus of Dolma (Soldaini et al., 2024).
@@
| Belebele | 0.7093 | 0.5633 | 0.7684 | 0.5705 | 0.5533 | 0.3028 | 0.2755 | 0.4457 | 0.3627 | 0.5057 | |
| m-MMLU | 0.5582 | 0.4734 | 0.6384 | 0.4720 | 0.4664 | 0.3260 | 0.2807 | 0.4207 | 0.3390 | 0.4416 | |
| m-ARC | 0.4779 | 0.4220 | 0.5054 | 0.3941 | 0.3892 | 0.3174 | 0.2970 | 0.3662 | 0.2731 | 0.3825 | |
| FLORES | 0.5088 | 0.3815 | 0.4110 | 0.3963 | 0.3939 | 0.0866 | 0.1946 | 0.2642 | 0.0413 | 0.2976 | |
| 0.7194 | 0.5872 | 0.7725 | 0.6538 | 0.6520 | 0.2464 | 0.3579 | 0.4716 | 0.1965 | 0.5175 | ||
| Bible | 0.3568 | 0.2152 | 0.3169 | 0.2103 | 0.2026 | 0.1246 | 0.0908 | 0.1198 | 0.0121 | 0.1832 | |
| 0.6076 | 0.4021 | 0.6599 | 0.4212 | 0.4190 | 0.2724 | 0.2357 | 0.2606 | 0.0319 | 0.3678 |
4.3 Benchmarks
Among the existing evaluation benchmarks in Table 5 from Appendix A.2, we chose Belebele (Bandarkar et al., 2024), m-ARC (Lai et al., 2023b), and m-MMLU (Lai et al., 2023b), which support the highest number of high-, medium-, and low-resource languages and are directly related to natural understanding tasks, which is the primary focus of this paper.
We use the entire test set for each of these benchmarks (§A.4 for details) for more details) to evaluate the models, except in one case. For Llama 3.1 70B, we use the first 500 questions of m-MMLU instead of the whole set due to resource constraints. Since the selected LLMs used in our experiment are not instruction-tuned, we use 5-shot in-context learning with the lm-evaluation-harness framework, employing log-likelihood-based multiple-choice scoring. Other settings, such as prompt templates, are configured according to the framework’s default (Gao et al., 2023; Biderman et al., 2024).
4.4 Evaluation Measures
We calculate the Pearson correlation coefficient to assess the strength of the correlation between MEXA and downstream performance on our evaluation benchmarks. This coefficient is a statistical measure of the strength and direction of the linear relationship between two variables. A high value would indicate that MEXA provides a reliable assessment of multilingual capabilities in English-centric LLMs.
5 Results
Table 1 presents the downstream performance of the selected models across three benchmarks, along with MEXA scores from two parallel datasets. Notably, among models with parameter sizes ranging from 7 to 9 billion, both Gemma 2 and Llama 3.1 outperform the other LLMs in terms of non-English downstream performance and MEXA scores. The Llama 3.1 and Llama 3 models exhibit similar alignment and downstream task performance, and both represent substantial advancements compared to Llama 2. Moreover, results for the Llama 3.1-70B model indicate that scaling can significantly enhance alignment when compared to its smaller version. Interestingly, while Mistral achieves comparable results to Gemma 1 on English benchmarks, it demonstrates inferior alignment, which likely accounts for its reduced performance on non-English tasks. Furthermore, the Llama 2 model achieves higher MEXA scores than OLMo, indicating better alignment. However, due to Llama 2’s weaker performance on English tasks, it fails to transfer this alignment effectively, leading to comparable non-English performance between Llama 2 and OLMo. This observation is further explored in Section 5.2, where we normalize the expected performance based on the pivot language, namely English.
@@ 5.1 MEXA Correlation Analysis
| Avg. across models | ||
| FLORES weighted average | Belebele | 0.8994 |
| Belebele | 0.9098 | |
| m-MMLU | 0.9513 | |
| m-MMLU | 0.9188 | |
| m-ARC | 0.9393 | |
| m-ARC | 0.8856 | |
| last token | Belebele | 0.9168 |
| Belebele | 0.9058 | |
| m-MMLU | 0.9545 | |
| m-MMLU | 0.9134 | |
| m-ARC | 0.9195 | |
| m-ARC | 0.8685 | |
| Bible weighted average | Belebele | 0.8496 |
| Belebele | 0.8811 | |
| m-MMLU | 0.8823 | |
| m-MMLU | 0.8210 | |
| m-ARC | 0.9018 | |
| m-ARC | 0.8354 | |
| last token | Belebele | 0.8147 |
| Belebele | 0.8070 | |
| m-MMLU | 0.7572 | |
| m-MMLU | 0.6998 | |
| m-ARC | 0.7469 | |
| m-ARC | 0.6885 |
We compute sentence embeddings for the selected models using two methods: weighted average based on token positions and last token (see §3.1). We apply mean and max pooling on the MEXA alignment scores across all model layers to derive a single score for each language. In Table 2 (refer to Table 6 for the detailed table), we report the correlation between the MEXA scores (computed using both mean- and max-pooling, for the two embedding methods) and task performances. Across all settings, the best overall result (higher correlation) is achieved when embeddings are computed using the weighted average, with mean pooling as the pooling method. We adopt this configuration as the default setting for MEXA.
FLORES vs Bible. In the default setting, the average Pearson correlation coefficient for the FLORES parallel dataset across different tasks is 0.9300, and while for the Bible parallel dataset, it is 0.8779. The reason the Bible scores are generally lower than FLORES is that FLORES data is cleaner and more aligned with modern, standardized texts, whereas the Bible data is older and more specialized. For example, for some languages, the orthography of Bible texts no longer matches today’s orthography. In the Bible, Arabic often includes diacritics, which are typically omitted in modern writing and tasks, making the text less familiar to models trained on contemporary data. Additionally, although the Bible dataset has been made parallel, sentence alignment can still be inconsistent due to translation nuances. In contrast, FLORES is carefully curated to ensure high-quality, sentence-level parallelism across languages for machine translation tasks.
5.2 Downstream Performance Estimation
A full score on Pearson correlation (i.e., ) indicates that a linear equation perfectly describes the relationship between MEXA and the evaluation benchmarks, with all data points lying on a line. Given the high correlation values shown in Table 6, it is reasonable to conclude that we can fit a line that closely approximates this linear relationship. This line converts the MEXA scores back to downstream task performances. We employed a linear model to predict this line by minimizing the residual sum of squares between the MEXA scores (multiplied by the performance on the English task) and the task performances. We needed to adjust the MEXA scores for this purpose, as the MEXA score for language indicates how well is aligned with English but does not reflect the estimated task performance of the model for language . Of course, this does not change the value of the correlation coefficient, as it is unaffected by linear transformations. The three tasks considered in this paper involve multiple-choice questions with four possible answers for each question, resulting in a chance of being randomly correct of . However, the minimum score for MEXA scores is 0. Thus, the ideal slope for the line would be with an intercept of (X-axis: adjusted MEXA scores, Y-axis: task performance). In Figure 2, we plot this relationship for Llama 3.1-8B using the Bible and FLORES parallel datasets for Belebele and m-ARC. We chose max pooling for Belebele and mean pooling for m-ARC, since these pooling methods yield a stronger correlation (see §5.1). The pairs of (slope, intercept) from left to right in the Figure 2 are: (0.6804, 0.2477), (0.6103, 0.1838), (0.6340, 0.3408), and (0.5726, 0.2423). With data points from both high-resource and low-resource languages, this line can be calculated; otherwise, the ideal line may be used as a reference.
5.3 MEXA vs Absolute Cosine Similarity
We compare MEXA with the use of absolute cosine similarities. We used parallel data from FLORES and downstream task data from the Belebele benchmark, focusing on 116 common labels. For each non-English language, we computed the average absolute cosine similarity for parallel sentences with English, and for non-parallel sentences with English. Following the setup by Li et al. (2025), which uses absolute cosine similarity values to predict language performance and ranking, we computed sentence embeddings using the last-token method and applied mean pooling over layers {5, 10, 15, 20, 25}.
To evaluate the correlation of each method with downstream task performance, we report results using the Gemma 1 and Llama 1 7B models. Table 3 summarizes these results. For both models, MEXA consistently achieves a higher correlation with downstream performance compared to the absolute cosine similarity of parallel sentences. Moreover, the correlation between cosine similarity for parallel and non-parallel sentences is notably high, suggesting that absolute values may be less discriminative across sentence types. This discrepancy highlights a limitation of using absolute cosine similarity: for some languages, similarity scores may remain high even for non-parallel sentences. Conversely, a low overall similarity score does not necessarily imply weak alignment, as parallel sentence scores may still significantly exceed non-parallel ones. In contrast, MEXA offers a more robust and comparative measure across languages.
| Gemma 1 7B | Llama 1 7B | |
| MEXA Belebele | 0.9260 | 0.8365 |
| AC-P Belebele | 0.7651 | 0.6473 |
| AC-P AC-NP | 0.9232 | 0.9064 |
5.4 Visualization of Layers
In Figure 3, we show the results of applying MEXA to 20 pairs of language_script from the FLORES parallel dataset for Llama 1-7B and Llama 3.1-8B across all 32 layers. We selected these languages from different families, writing systems, and both high- and low-resource categories. The embeddings are computed using a weighted average based on token positions. Figure 3 shows that high-resource languages (with more prevalence on the web; see §A.1) achieve higher alignment scores across different layers, while low-resource languages achieve lower scores. In the initial layers, embeddings are more in-language, resulting in lower alignment scores. As embeddings progress to the mid-layers, they become more aligned with the dominant language of the LLM, i.e., English.
MEXA is comparable between models as long as the same parallel dataset and setting is used to obtain the MEXA scores. Figure 3 shows that in many languages, particularly high-resource languages, Llama 3.1 achieves a significantly higher alignment score than its predecessor, Llama 1. Although Llama 3.1 leads to better alignment scores with English for medium and low-resource languages, there is still room for improvement. Comparing Arabic (arb_Arab) with its romanized version (arb_Latn), we see that both Llama 1 and Llama 3.1 models perform better in the native script than in the Latin script, even though Llama 1’s tokenizer for Arabic is essentially a character-based tokenizer. In general, for very low-resource languages, those in Latin script tend to have higher alignment scores, likely because the tokenization is more favorable for Latin characters.
In Figure 4, we display the t-SNE (Van der Maaten and Hinton, 2008) plots of the embeddings of Figure 3 from 3 different layers of Llama 3.1: embedding layer 0, mid-layer 13, and last layer 32. We assign a different color to each language. For layers 0 and 32, the embeddings are more language-specific, while in the mid-layer, they become more language-neutral. Languages that maintain their language-specific embeddings in the mid-layer are clustered separately and, notably, correspond to the very low-resource languages that receive the lowest alignment scores from MEXA.
6 Conclusion
We introduce MEXA, a method for assessing the multilingual capabilities of English-centric large language models (LLMs). MEXA builds on the observation that English-centric LLMs semantically use English as a kind of pivot language in their intermediate layers. MEXA computes the alignment between non-English languages and English using parallel sentences, estimating the transfer of language understanding capabilities from English to other languages through this alignment. This metric can be useful in estimating task performance, provided we know the English performance in the task and the alignment score between languages derived from a parallel dataset. Through different studies with two parallel datasets (FLORES-200 and the Bible), different LLMs including the Llama family, Gemma family, Mistral, and OLMo, and three downstream tasks (Belebele, m-MMLU, and m-ARC), we demonstrated that MEXA provides a reliable estimation of multilingual performance. For MEXA score calculations, multiple design analyses are conducted to explore the impact of token-level pooling for embeddings and layer-level pooling in computing alignment scores. While MEXA shows high correlation across most configurations, a weighted average of tokens combined with mean pooling delivers the best results. The results reveal a promising average value for the Pearson correlation coefficient of 0.90 with established downstream tasks across nine models and two parallel datasets. Overall, MEXA proves to be a valuable method for practitioners aiming to assess the multilingual capabilities of English-centric LLMs, easing future efforts to expand these models to a wider range of underrepresented languages.
Limitations
We are aware of four main limitations of our work.
First, The scope of this paper is limited to non-generative tasks. Generation is generally more challenging than understanding, and it is unsurprising that these models for many languages may struggle to generate content in their language. While NLP has advanced toward generative capabilities, a significant portion of evaluation still focuses on non-generative tasks (e.g., Hugging Face Leaderboards)222hf.co/spaces/open-llm-leaderboard/open_llm_leaderboard due to their convenience in multiple-choice question evaluation and standardized metrics. Assessing generated output remains challenging, even in English benchmarks. For example, model-based approaches (e.g., “LLM-as-a-judge” (Zheng et al., 2023)) require an LLM fully competent in the target language—a capability that is both questionable and the focus of our evaluation.
Second, MEXA provides a method of evaluation for open science, and only model weights are needed. Although, developers of closed-source models could use MEXA under the hood and report their multilingual results to provide insight of their model’s multilingual capabilities. We target the widely used settings where the LLM follows a decoder-only transformer architecture. For other architectures, as long as we can extract the embedding given a sentence for intermediate layers, MEXA can be calculated.
Third, We present a selection of tasks for multilingual evaluation in Table 5. As shown, for non-generative tasks, only a few benchmarks support a high number of languages, including low-resource ones. Benchmarks limited to around 10 languages, which mostly support high-resource languages, would not support our claims, as MEXA would achieve high results for all of them. Belebele includes the highest number of languages (except sequence labeling tasks), making it an ideal task to evaluate MEXA. Both m-MMLU and m-ARC are the next highest covered languages for non-generative tasks. However, since they are machine-translated tasks, they are not ideal and may bias some results for low-resource languages (or, more accurately, when the machine translation is poor). Yet, these translated versions are representative of the current state of automatic evaluation, as seen in multilingual leaderboards.333hf.co/spaces/uonlp/open_multilingual_llm_leaderboard
Fourth, MEXA provides a rough estimate of the multilingual capabilities of pre-trained English-centric LLMs. Different tasks offer diverse perspectives on the abilities of LLMs, and MEXA cannot replace all of them. Our goal is to highlight the multilingual potential of English-centric LLMs and propose a simple way to evaluate them. We hope this encourages the development of more multilingual LLMs, even though they are likely to contain large shares of English data. Additionally, it is important to note that answers across languages do not always need to be fully aligned (Naous et al., 2024), and for such cases, language- and culture-specific evaluation benchmarks should be developed.
Acknowledgments
Appendix A Appendix
A.1 Distribution of Pre-training Data in LLMs
The distribution of languages in the training data of state-of-the-art LLMs is rarely fully documented. Llama 2 (Touvron et al., 2023b) is a counter-example and its authors have disclosed the language distribution use in pre-training. Their analysis uses the FastText (Bojanowski et al., 2017) language identification tool and a threshold of for the language detection. We reproduce Touvron et al. (2023b, Table 10), which lists 27 languages with percentages greater than 0.005% in the Llama 2 pre-training data, in Table 4. English, with 89.70%, constitutes the vast majority of the training data.
The distribution of data in the training of English-centric LLMs is not the same as on the web, but it does have some correlation. The amount of English in LLM pre-training data is significantly larger than for other languages. This is also observable for GPT-3 (Brown et al., 2020b), where more than 92% of the training texts was in English (Brown et al., 2020a). The rest of the top languages in the data of such models are mostly high-resource languages, which have the most available data on the web (top 36 languages). However, in some models, this could be adjusted by design, for example, to make writing systems with non-Latin languages less prominent (as seen in Llama 2). This weakens the correlation between LLMs’ pre-training data and the web.
| Language | Script | Percent | Language | Script | Percent |
| English (eng) | Latn | 89.70% | Ukrainian (ukr) | Cyrl | 0.07% |
| Unknown (unk) | - | 8.38% | Korean (kor) | Hang | 0.06% |
| German (deu) | Latn | 0.17% | Catalan (cat) | Latn | 0.04% |
| French (fra) | Latn | 0.16% | Serbian (srp) | Cyrl/Latn | 0.04% |
| Swedish (swe) | Latn | 0.15% | Indonesian (ind) | Latn | 0.03% |
| Chinese (zho) | Hans/Hant | 0.13% | Czech (ces) | Latn | 0.03% |
| Spanish (spa) | Latn | 0.13% | Finnish (fin) | Latn | 0.03% |
| Russian (rus) | Cyrl | 0.13% | Hungarian (hun) | Latn | 0.03% |
| Dutch (nld) | Latn | 0.12% | Norwegian (nor) | Latn | 0.03% |
| Italian (ita) | Latn | 0.11% | Romanian (ron) | Latn | 0.03% |
| Japanese (jpn) | Jpan | 0.10% | Bulgarian (bul) | Cyrl | 0.02% |
| Polish (pol) | Latn | 0.09% | Danish (dan) | Latn | 0.02% |
| Portuguese (por) | Latn | 0.09% | Slovenian (slv) | Latn | 0.01% |
| Vietnamese (vie) | Latn | 0.08% | Croatian (hrv) | Latn | 0.01% |
A.2 Multilingual Evaluation Benchmarks
| Dataset | Task | # L |
| XNLI (Conneau et al., 2018) | Natural Language Inference | 15 |
| IndicXNLI (Aggarwal et al., 2022) | Natural Language Inference | 11 |
| AfriXNLI (Adelani et al., 2024) | Natural Language Inference | 15 |
| m_HellaSwag (Lai et al., 2023b) | Natural Language Inference | 31 |
| PAWS-X (Yang et al., 2019) | Paraphrase Identification | 7 |
| XCOPA (Ponti et al., 2020) | Commonsense Reasoning | 11 |
| XStoryCloze (Lin et al., 2022) | Commonsense Reasoning | 11 |
| m-ARC (Lai et al., 2023b) | Common Sense Reasoning | 31 |
| TyDiQA (Clark et al., 2020) | Question Answering | 11 |
| MLQA (Lewis et al., 2020) | Question Answering | 7 |
| XQuAD (Artetxe et al., 2020) | Question Answering | 11 |
| IndicQA (Doddapaneni et al., 2023) | Question Answering | 10 |
| AfriQA (Ogundepo et al., 2023) | Question Answering | 10 |
| m_TruthfulQA (Lai et al., 2023b) | MC General Question Answering | 31 |
| UDPOS 2.7 (de Marneffe et al., 2021) | Part of Speech Tagging | 104 |
| WikiANN (Pan et al., 2017) | Name Entity Recognition | 282 |
| XLSum (Hasan et al., 2021) | Summarization | 44 |
| WikiLingua (Ladhak et al., 2020) | Summarization | 18 |
| Belebele (Bandarkar et al., 2024) | MC Reading Comprehension | 115 |
| AfriMMLU (Adelani et al., 2024) | MC Knowledge Question Answering | 17 |
| m-MMLU (Lai et al., 2023b) | MC Knowledge Question Answering | 31 |
| MMMLU (OpenAI, 2024) | MC Knowledge Question Answering | 15 |
| M3Exam (Zhang et al., 2023a) | MC Multimodal Question Answering | 9 |
A.3 Semantic Similarity in Multilingual Embeddings
There are other ways to compute similarity between languages, such as Representational Similarity Analysis (RSA) (Chrupała and Alishahi, 2019) and Central Kernel Alignment (CKA) (Kornblith et al., 2019). RSA involves first computing the cosine similarity for sentence embeddings within each language, then correlating these in-language similarities with those in other languages. CKA, another metric, is adopted by Conneau et al. (2020b) and Muller et al. (2021). Conneau et al. (2020b) show that the CKA similarity is highly correlated with sentence retrieval scores for four languages. In this paper, our focus is not on finding different ways to calculate similarity between languages, but on how helpful a properly defined alignment score can be in estimating the multilingual capabilities of LLMs across multiple languages.
A.4 Benchmark Details
Belebele is a multiple-choice reading comprehension task designed to assess language models across a range of high-, medium-, and low-resource languages. Each question in the dataset is paired with four possible answers and linked to a brief passage from the FLORES-200 dataset (NLLB Team et al., 2022). The human annotation process was carefully curated to generate questions that effectively differentiate between various levels of language comprehension, supported by rigorous quality checks. Belebele supports 122 distinct labels (language-script combinations) corresponding to 115 distinct languages. However, FLORES-200 does not support 5 of these labels, corresponding to Romanized versions of 5 Indic languages. Therefore, we conducted our analysis between the FLORES-200 and Belebele benchmarks on 117 common labels. Additionally, there are 102 common labels between the Bible parallel data and the Belebele benchmark.
Both ARC Challenge (Clark et al., 2018) and MMLU (Hendrycks et al., 2021) are also set up as multiple-choice question-answering tasks, but they focus on different types of knowledge and reasoning skills. ARC Challenge is classified as a common-sense reasoning task, consisting of grade-school level science questions, while MMLU consists of questions across a wide range of fields, including the humanities, social sciences, and more. Lai et al. (2023b) used GPT-3.5-turbo (OpenAI, 2022) and a translation prompt to translate examples from both datasets and create m-ARC and m-MMLU in 31 languages (excluding English). Later, m-MMLU was expanded to also include Icelandic (isl_Latn) and Norwegian (nob_Latn). The Icelandic portion was translated using the Mideind.is, while Norwegian was generated with DeepL.com.444hf.co/datasets/alexandrainst/m_mmlu m-MMLU consists of 277 questions in its training set, 13,258 in the test set and 1,433 in the validation set. m-ARC consists of 1,116 questions in the training set, 1,169 in the test set, and 298 in the validation set.